a number,twice that number,and 1/5 of that number are add,the result is 16?
Found 3 solutions by nyc_function, mananth, stanbon:
Answer by nyc_function(2741)

(1/5)(x) can be written as x/5. They both mean the same thing.
x + 2x + (x/5) = 16
I gave you the equation. Can you finish?

Answer by mananth(16949)

let the number be x
twice the number = 2x
1/5 of the number = x/5
..
x+2x+x/5 = 16
LCM =5
5x+10x+x= 16*5
16x=80
x=5 the number

Answer by stanbon(75887)

a number,twice that number,and 1/5 of that number are add,the result is 16?
--------------
Let the number be "x":
Equation:
x + 2x + (1/5)x = 16
3 1/5 x = 16
---
(16/5)x = 16
(1/5)x = 16
x = 80
===================
Cheers,
Stan H.